Although appealing to more refined tastes, art as a collectible has not always performed so profitably. During 2015, an auction house sold a painting for a price of $1,190,000. Unfortunately for the previous owner, he had purchased it three years earlier at a price of $1,770,000. What was his annual rate of return on this painting?
Answer #1

The return has been negative in this case because he had purchased for 1,770,000 and currently it is being sold at 1,190,000. So, the value has decreased over a period of 3 years. The decrease in value can be calculate as:

1,190,000 = 1,770,000 * (1+r)^3

(1+r)^3 = 0.67

1+r = 0.87

r = -0.123

It means, the price has fallen by 12.3% every year for over a period of 3 years.